OpenStudy (anonymous):

**MEDAL/FAN A study of five hundred adults found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 14 hours. The population standard deviation is 3 hours. What is the margin of error for a 95% confidence interval? 0.134 0.220 0.263 0.313

OpenStudy (anonymous):

@ganeshie8 @pooja195 @Michele_Laino

OpenStudy (michele_laino):

please wait, I'm pondering...

OpenStudy (anonymous):

Yes okay

OpenStudy (michele_laino):

a 95% confidence interval, means an error of: \[\Large \frac{{1.96\sigma }}{{\sqrt {500} }} = \frac{{1.96 \times 3}}{{\sqrt {500} }} = ...?\]

OpenStudy (michele_laino):

I got that number: 1.96, using my tables of the "erf" function

OpenStudy (michele_laino):

N=500 is the size of your sample

OpenStudy (anonymous):

5.88 ---- 22.36067977

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

0.2629615942

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Can u help me with this? A coach wants to know if football players who warm up before a game will score more points than those who don't warm up. He will break the team up into two groups and make sure each player gets equal time on the field. He will then count how many points each player makes during the game. What statistic should he study? The mean number of points earned by each player The mean number of points earned by each group The standard deviation of the number of points earned by each player The standard deviation of the number of points earned by each group

OpenStudy (michele_laino):

I think the last option

OpenStudy (michele_laino):

since the coach is searching for an average value for player

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

Can you help me with one last question? Crunchy Potato Chip Company has two different manufacturing plants. Company officials want to test whether each plant fills the bags with the same number of ounces. A random sample of potato chip bags from plant A had a mean of 24.2 ounces in each bag. A random sample from plant B had a mean of 22.2 ounces. They randomized the data over 100 trials and the difference of means for each trial is shown in the dot plot below. What can Crunchy Potato Chip Company conclude from this study?

OpenStudy (anonymous):

Options: The difference is not significant because a difference of 2.0 is very likely. The difference is not significant because a difference of 2.0 is not very likely. The difference is significant because a difference of 2.0 is not very likely. The difference is significant because a difference of 2.0 is very likely.

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (michele_laino):

I think the second option, since we have no outcomes whose difference is 2 ounces

OpenStudy (anonymous):

Thank you :)

OpenStudy (michele_laino):

please wait

OpenStudy (michele_laino):

I confirm my answer

OpenStudy (anonymous):

Okay

OpenStudy (michele_laino):

If your problem is searching for a difference 2, in absolute value, then we have to consider also the result -2.0

OpenStudy (michele_laino):

in that case the right option is option A, since outcomes for a difference about -2.0 give a low probability

OpenStudy (anonymous):

Oh okay! Thank you very much :)

OpenStudy (michele_laino):

thanks! :)